Determine the work done by the force F(xy+3z)i^+(2y^2 -x^2)j+(z-2y)k in taking a particle from  x = 0 to  x =1 along a curve defined by the equations: x^2= 2y ; 2x^3= 3z
1
Expert's answer
2020-04-06T08:56:14-0400
"\frac{dy}{dx}=x,\frac{dz}{dx}=2x^2"

"W=\int_0^1((xy+3z)+(2y^2 -x^2)x+(z-2y)2x^2)dx"

"W=\int_0^1Idx"

"(xy+3z)=x\frac{x^2}{2}+2x^3=\frac{5x^3}{2}"

"(2y^2 -x^2)x=\left(2\left(\frac{x^2}{2}\right)^2 -x^2\right)x=\left(\frac{x^5}{2} -x^3\right)"

"(z-2y)2x^2=\left(\frac{2x^3}{3}-x^2\right)2x^2=\left(\frac{4x^5}{3}-2x^4\right)"

"I=\frac{5x^3}{2}+\left(\frac{x^5}{2} -x^3\right)+\left(\frac{4x^5}{3}-2x^4\right)=\left(\frac{11x^5}{6}+\frac{3x^3}{2}-2x^4\right)"

"W=\int_0^1\left(\frac{11x^5}{6}+\frac{3x^3}{2}-2x^4\right)dx"

"W=\left(\frac{11}{6(6)}+\frac{3}{2(4)}-2\frac{1}{5}\right)=\frac{101}{360}\approx 0.28"